A study by Siepielski and Benkman demonstrated phenotypic selection on pinecones by granivorous (seed-­‐eating) birds. Specifically, they found that interactions between the birds and pine trees span from antagonistic (birds are predators and confer a fitness disadvantage to the pines) to mutualistic (birds still eat seeds, but they cache what they cannot eat and confer a dispersal benefit to the pines) depending on the number of seeds produced in a population. Assume seed production is determined by two alleles that demonstrate incomplete dominance (i.e. additive genetic variance): high (F) and low (f) seed production. Please answer the following questions:

- In populations with high seed production, relative fitness of the pines was found to be approximately FF = 1.00, Ff = 0.63, and ff = 0.13. Assuming an initial allele frequency of P(F) = 0.5, an effective population size of 100, and 10 independent populations:

  1. If F is dominant, how many generations (approximately) does it take for F to become fixed?
  2. If F is recessive, how many generations (approximately) does it take to become fixed?
  3. Why do we see this change?
Answer #1

A) solution through allele frequency method

  • FF Ff ff
  • 1.00 0.63    0.13
  • 2 (F) 0.63(F)  0.63(f) 0.26(f)
  • 2+0.63 0.26+0.63
  • 2.63 0.89

high seed production. low seed production.

2.63+0.89 = 3.52  

2.63/3.52= 0.74 high seed production

0.89/3.52= 0.25 low seed production

So if F is dominant it takes 0.74 high seed production

B) F is recessive it takes 0.25 low seed production